the side of a square equals the length of a rectangle. the width of the rectangle is 4 meters longer than its length. The sum of the areas of the square and the rectangle is 240 square centimeters. find the side of the square.
Answer by Boreal(15235)

side of square=s
dimensions of rectangle are s and s+4
Area of square is s^2
Area of rectangle is s^2+4s, product of l*w
They add to 240 cm^2
Therefore, s^2+s^2+4s=240
2s^2+4s-240=0
2(s^2+2s-120)=0
(s+12)(s-10)=0
s=10,-12
use positive root only
square area is 10*10=100
rectangle area is 10*14=140
They add to 240 sq cm
side is 10 cm